OpenStudy (anonymous):

A circular camera lens has an area of about 19(PIE SIGN) square millimeters. What is the diameter of the camera lens?

OpenStudy (anonymous):

Help please

OpenStudy (anonymous):

helllp

OpenStudy (anonymous):

Hi :) You know how to get the circumference of a circle when given its diameter?

OpenStudy (anonymous):

Or, sorry, the area of a circle, given its diameter.

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Well, when you have a diameter, how do you get the radius?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

I forgot these things cuz my class was thought advanced things

OpenStudy (anonymous):

Well, the way to get an area of a circle, given its diameter, first you get the radius of the circle, which is just half the diameter \(\large r=\frac12\times d\). Then the area of a circle is pi times the radius, times the radius (again) \[\Large A=\pi \times r\times r\]

OpenStudy (anonymous):

how do you get the diameter of a circle with a GIVEN area

OpenStudy (anonymous):

Just reverse the process ^.^

OpenStudy (anonymous):

what do you mean?!

OpenStudy (anonymous):

I mean... \[\Large A=\pi r^2\], put the area in, and then solve for r ^.^

OpenStudy (anonymous):

but how can I get the radius if I don't have the diameter!!!

OpenStudy (anonymous):

You can also get the radius from the area, silly :) Using THAT formula \(\huge A=\pi r^2\)

OpenStudy (anonymous):

? I don't understand :\

OpenStudy (anonymous):

What is A?

OpenStudy (anonymous):

A=59.69

OpenStudy (anonymous):

Oh, don't put it like that, I suggest you leave it as \(\large 19\pi\) for now :)

OpenStudy (anonymous):

ok. so A=\[19\]

OpenStudy (anonymous):

MY BAD..A=19PIE

OpenStudy (anonymous):

So.. \[\huge 19\pi=\pi r^2\] NOW solve for the radius, r.

OpenStudy (anonymous):

UHY!I DONT UNDERSAND..I HAVE TO GO NOW..

OpenStudy (anonymous):

Okay :) Bye, and good luck ^.^